OpenStudy (anonymous):

Solve to find X. >>> x+2=sqrt(7+2x)

OpenStudy (anonymous):

x+2=sqrt(7+2x) square both sides: (x + 2)^2 = 7 + 2x x^2 + 4x + 4 = 7 + 2x x^2 + 2x - 3 = 0 do you know what to do from here?

OpenStudy (anonymous):

ummmm.... not really. could you explain then i could try ..

OpenStudy (anonymous):

subtract 2x and -3?

OpenStudy (anonymous):

ok, factor that left side of x^2 + 2x - 3 = 0:

OpenStudy (anonymous):

(x+3)(x-1)

OpenStudy (anonymous):

yes... now the solution is obviously x = -3 or x = 1. But remember to always check these solutions by sticking your solutions back into the ORIGINAL equation. What's the solution?

OpenStudy (anonymous):

okkkk OMG thank youuu soo much.

OpenStudy (anonymous):

way to go! you're welcome.